Circular No. 7433

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMETS C/2000 J5, C/2000 K3, C/2000 K4, C/2000 K5 D. A. Biesecker, Emergent Information Technologies, Inc., and God-

dard Space Flight Center, reports observations of four additional Kreutz sungrazing comets detected by SOHO with the LASCO C2 telescope. The comets were found via the SOHO website by M. Oates (C/2000 J5, C/2000 K4, C/2000 K5) and by M. Boschat (C/2000 K3, C/2000 K5). All but C/2000 K5 showed a tail, those of C/2000 K3 and C/2000 K4 being noted as short. Full astrometry (including reductions by B. G. Marsden) and orbital elements appear on the MPECs indicated below.

2000 UT  α_{2000}  δ_{2000}  MPEC  Comet
May 11.810 3^{h}16.^{m}5 +16^{o}31′ 2000-K34  C/2000 J5
19.893 3 49.9 +18 26  2000-K35  C/2000 K3
22.576 4 00.4 +18 59  2000-K36  C/2000 K4
25.979 4 13.3 +19 32  2000-K37  C/2000 K5

SUPERNOVA 2000cf IN MCG +11-19-25 K. Krisciunas and A. Rest, University of Washington; and R. McMil-

lan, Apache Point Observatory (APO), report preliminary photometry for SN 2000cf, obtained with the APO 3.5-m telescope: On May 11.18 UT, V = 17.09, B − V = +0.08, V − R = +0.10, V − I = −0.11 (uncertainties ± 0.03 mag). On May 20.17, V = 17.44, B − V = +0.48, V − R = +0.10, V − I = −0.05 (uncertainties ± 0.04 mag). Near-infrared photometry on May 11.27 yields H = 16.79 ± 0.14 (affected by a rather serious gradient in the underlying galaxy light).

DM GEMINORUM Y. Lipkin, E. Leibowitz, and A. Retter, Wise Observatory, report:

“We performed CCD photometry of DM Gem (N Gem 1903) with the Wise Observatory 1-m telescope between 1999 Nov. 11 and 2000 Mar. 30, and a power-spectrum analysis of the I light reveals a clear and significant signal of quasiperiodic oscillations. These are interpreted as a combination of two distinct periodicities of 0.12287 ± 0.0001 and 0.12266 ± 0.0001 day, each with semiamplitude ∼ 0.03 mag. A similar set of measurements was obtained 2 yr earlier, between 1997 Nov. 12 and 1998 Mar. 21, when the star was on average 0.125 mag brighter in I; no trace of any periodicity is apparent in that light curve.”

2000 May 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green